Solve each equation by first eliminating the fractions
1/2q - 6 = 1/5q
Solve each equation by first eliminating the decimal numbers
0.5b + 3.4 = 0.2b + 12.4
Solve each equation. If you feel proficient enough, try simplying the process
4x - 2 = -8 + 5x
Solve each equation. Identify each as a conditional equation, an inconsistent equation or an identity
3 - 3(5 - x) = 0
Solve each equation
-3 - 4(t - 5) = -2(t + 3) + 11

Answer by zeynep(43)

1/2q - 6 = 1/5q
let's get rid of those fractions by multiplying moth sides by the LCD of 10;
10.(1/2q - 6) =10.( 1/5q )
5q-60=2q add additive inverse of 2q to both sides;
5q-60-2q=2q-2q
3q-60=0 add additive inverse of -60 to both sides;
3q-60+60=0+60
3q=60 divide both sides by 3;
q=20
0.5b + 3.4 = 0.2b + 12.4
Let's multiply both sides by 10 to eliminate the decimals;
5b+34=2b+124
5b+34-2b=2b+124-2b
3b+34=124
3b+34-34=124-34
3b=90 divide both sides by 3;
b=30
4x - 2 = -8 + 5x
get the variables on the same side by adding additive inverse of 5x to both sides.
4x - 2-5x = -8 + 5x-5x
-x-2=-8 add additive inverse of -2 to oth sides to leave the variable, which is x, alone;
-x-2+2=-8+2
-x=-6 divide both sides by -1;
x=6

3 - 3(5 - x) = 0
3-15+3x=0
-12+3x=0
-12+3x+12=0+12
3x=12 divide both sides 3;
x=4
This equation is a conditional equation. A conditional equation is an equation that is not true for all real numbers but has at leat one real number solution.

-3 - 4(t - 5) = -2(t + 3) + 11
-3-4t+20=-2t-6+11
-4t+17=-2t+5
-4t+17+2t=-2t+5+2t
-2t+17=5
-2t+17-17=5-17
-2t=-12 divide both sides by -2
t=6